OpenStudy (anonymous):

Solve the equation below for x: log(5X)+log(x-1)=2 A. There is no solution. B. x = 5 C. x = 4; x = -4 D. x = 5/2

OpenStudy (slaaibak):

first combine them: log(5x(x-1)) = 2 10^2 = 5x(x-1) 5x^2 - 5x - 100 = 0 x^2 - x - 20 = 0 (x+4)(x-5) = 0 x=-4 or x=5 but you can't take the log of a negative number, so -4 doesn't work. so the answer is 5.